OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -11 or x + 5 greater than or equal to 14. A number line with a closed circle on -8, shading to the left, and an open circle on 9, shading to the right A number line with an open circle on -8, a closed circle on 9, and shading in between A number line with an open circle on -8, shading to the left, and a closed circle on 9, shading to the right A number line with a closed circle on -8, an open circle on 9, and shading in between

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@hartnn

OpenStudy (gorv):

x - 3 < -11 or x + 5 greater than or equal to 14. x-3<-11 or x+5>=14 x<-11+3 or x>=14-5 x<-8 or x>=9

OpenStudy (anonymous):

so which answer i am very confused

OpenStudy (gorv):

open circle come if equal sign is not there if equal to sign is given then shaded as given in x>=9

OpenStudy (gorv):

now tell me x<-8 will shaded???

OpenStudy (anonymous):

yess :) right

OpenStudy (gorv):

you got it ??

OpenStudy (anonymous):

soo then the answer is B then

OpenStudy (gorv):

x<-8 mean our ans is less than -8 ...that is -9 -10 -11 -12...... x>=9 means our ans is 9 10 11 12 13 .....